# Crypto Inverse-Power Options and Fractional Stochastic Volatility

### v2

@@ 1 Introduction
In the present paper we introduce a class of fractional stochastic-volatility models that are capable of reconciling various prevailing stylized facts in financial markets. Broadly speaking, our main objective is to design a general model framework that accommodates both flexible jump dynamics and certain path-dependence properties, and simultaneously achieve fast, semi-analytical valuation. The proposed models address the shortcomings of various existing ones that struggle to provide comprehensive modeling and calibration within a reasonable time frame, while they also maintain parametric parsimony by adding only one or two optional parameters, depending on the desired jump behavior. Following this model framework, we revisit the pricing–hedging problem of derivatives in the cryptocurrency market, where significant price and volatility jumps have been well-documented. Particular attention is drawn to inverse contracts, namely those that are settled in cryptos (    C    ) despite the underlying being valued in a fiat currency (e.g., USD (\$)). [^3] Notably, around 90% of actively traded crypto options currently belong to the inverse type, whose payoffs can be effectively approximated by those of direct options post-conversion, similar to how foreign-exchange (FX) options work.
@@ 1.1 Jumps in crypto prices and volatility
From a derivative pricing perspective, crypto (especially Bitcoin) price dynamics have been explored extensively in recent studies. To mention a few, [Madan et al., 2019] [43] considered calibrating a series of Markov models on Bitcoin options, concluding that models with infinitely active jumps and volatility clustering are generally favorable in terms of pricing errors. This aspect was also strengthened in [Hou et al., 2020] [28], which evidenced the importance of jumps in crypto prices and volatility, as well as their co-movements, by considering both statistical estimation and option pricing under two stochastic-volatility models with correlated jumps from [Duffie et al., 2000] [19] and [Bandi and Renò, 2016] [6], respectively. Significantly, [Hou et al., 2020] [28] has documented an inverse leverage effect, indicating positive serial correlations between crypto price and volatility jumps. Notably, while the diffusive leverage effect may remain negative or insignificant, the inverse leverage effect is particularly pronounced during periods of large fluctuations, which phenomenon is uniquely captured by jump models. We refer to [Huang et al., 2022] [29] for a comprehensive discussion based on analyzing crypto time series data. Besides, the papers [Jalan et al., 2021] [34] and [Hilliard and Ngo, 2022] [26] also stressed, respectively, the importance of stochastic volatility and price jumps by comparing the pricing performances of a regime-switching stochastic-volatility model and a jump-diffusion model with stochastic convenience yield to that of the Black–Scholes model. Furthermore, [Cao and Celik, 2021] [13] studied the valuation of Bitcoin options under an equilibrium setting with various sources of jump risks, again highlighting the undervaluation of deeply-in-the-money options by the Black–Scholes model.
@@ 2 Model framework
We assume that the crypto market evolves in continuous time $t\geq 0$ and is governed by a risk-adjusted stochastic basis $(\Omega,\mathcal{F},\mathsf{Q},\mathbb{F}\equiv\{\mathscr{F}_{t}:t\geq 0\})$. The market filtration $\mathbb{F}$ satisfies the usual conditions and contains all market information, and $\mathsf{Q}$ is some preselect pricing measure. All stochastic processes to appear are understood to be supported on this basis.
@@ 2.1 Fractional stochastic volatility with jumps
To establish stochastic volatility when jumps occur in both price and volatility dynamics, a powerful technique that avoids explicitly specifying an instantaneous variance process is to incorporate a time-change argument that works by replacing the calendar time $\imath\equiv(t)$ with a random, constantly increasing business time, denoted as $\mathcal{T}\equiv(\mathcal{T}_{t})$. Intuitively speaking, the business time serves to capture variations in trading activity, incorporate return autocorrelations, and create volatility clustering. Its theoretical validity is deep-rooted in the so-called “theorem of time change” by [Monroe, 1978] [44], which states that any semimartingale can be represented as a time-changed Brownian motion. In the same vein, the time change of a general semimartingale is legitimate provided that the sample paths of $\mathcal{T}$ are continuous $\mathsf{Q}$-a.s. With the monotonicity of $\mathcal{T}$, this implies the typical construction
@@
In general, since volatility is mean-reverting in the long run, a familiar model choice for $A$ in the presence of jumps is a Lévy-driven Ornstein–Uhlenbeck process. In the present context, we modify the linear structure with kernel convolution for short-range dependence, by adapting [Wang and Xia, 2022, Sect. 2.2] [53]. We remark that such modification readily leads to fractional Ornstein–Uhlenbeck processes discussed in [Wolpert and Taqqu, 2004] [55], and these processes, having many attractive analytical properties, belong to the broader class of Lévy-driven Volterra processes detailed in [Barndorff-Nielsen et al., 2014] [7].
To be precise, let $Y\equiv(Y_{t})$ be a square-integrable Lévy subordinator (nondeterministic by default), with (Lévy) characteristics $(\mu_{Y},0,\nu_{Y})$ under $\mathsf{Q}$, where $\mu_{Y}\geq 0$ and the Lévy measure $\nu_{Y}$ is non-atomic and corresponds to a Poisson random measure $N_{Y}$ defined on $(\mathbb{R}_{++},\mathbb{R})$, such that $Y_{1}$ admits the characteristic exponent
@@
where $\kappa>0$ and $d\in(1/2,1)$ are a mean-reversion parameter and a fraction parameter. The fractional counterpart of $Y$ is constructed as the following $h$-convoluted stochastic integral:
@@
A notably simple choice of $h$ is the exponential-Riemann–Liouville product kernel (see (4.1.1)), $h(t,s)\equiv h(t-s)=e^{-\kappa(t-s)}(t-s)^{d-1}/\mathrm{\Gamma}(d)$, where $\mathrm{\Gamma}(\cdot)$ is the usual Gamma function and which is widely adopted in the rough volatility literature due to its structural simplicity, including in constructing the rough Heston model (see, e.g., [El Euch and Rosenbaum, 2019] [21]) and fractional Ornstein–Uhlenbeck processes in [Wolpert and Taqqu, 2004] [55]. We will consider other kernel forms as appearing in [Wang and Xia, 2022, Sect. 2.2] [53] later in Section 4.1, and some other special forms of $h$ can be found in [Barndorff-Nielsen et al., 2014, Sect. 3.1] [7].
@@ 2.2 Time-changed processes and crypto prices
According to (2.1.1), the business time incorporating fractional stochastic volatility is constructed as the time integral of $A$ in (2.1.5), and a major portion of shocks to the crypto price are governed by some $\mathcal{T}$-changed Lévy process, denoted $X\equiv(X_{t})$. Here, $X$ is a real-valued Lévy process with $\mathsf{Q}$-characteristics $(\mu_{X},\sigma_{X},\nu_{X})$, admitting the characteristic exponent
@@
Likewise, we have assumed that $X_{1}$ has finite moments. Also, to capture continual price fluctuations, it is reasonable to assume that either $\sigma_{X}>0$ or $\nu_{X}$ is infinite. The process $X$ is a base process describing price innovations without volatility clustering, which can already have jumps of flexible activity and scales. With stochastic volatility, the crypto price, henceforth the crypto/USD conversion rate, is formally given by
@@
where $S_{0}$ is the spot rate and $\rho\in\mathbb{R}$ is a leverage parameter channeling the (direct or inverse) leverage effect between prices and volatility. The existence of the (positive) quantities $\phi_{X_{1}}(-{\rm i})$ and $\phi_{Y_{1}}(-{\rm i}\rho)$ is the same as the requirement that $\mathsf{E}[e^{X_{1}}]<\infty$ and $\mathsf{E}[e^{\rho^{+}Y_{1}}]<\infty$. Together, these quantities act as exponential compensators to guarantee the martingale property of $S$ under $\mathsf{Q}$.
We remark that the price model (2.2.1) can be seen as a generalized version of the well-known Barndorff-Nielsen–Shephard model ([Barndorff-Nielsen and Shephard, 2001] [8]) in which $X$ is particularly a Brownian motion and $A$ is an ordinary Ornstein–Uhlenbeck process, i.e., one with $h(t,s)=e^{-\kappa(t-s)}$, $s\in[0,t)$. An extension where $X$ is a general Lévy process has been proposed and studied by [Yamazaki, 2016] [60], from which the key difference in the present framework mainly lies in the (singular) kernel function $h$.
To enable dynamic valuation of crypto derivatives, we need the conditional distribution of the log-price $\log S_{t}|\mathscr{F}_{t_{0}}$, for generic time points $0\leq t_{0}<t\leq T$, where $T>0$ is a predetermined expiry date. In the case with stochastic volatility, dynamic valuation is achievable provided that one has access to corresponding volatility derivatives, such as a variance swap which entitles the long party to receive a floating leg equal to the realized variance taken over the swap’s life at the expiry $T>0$ in exchange for a preset fixed payment such that the swap has zero value at the time of issuance. We refer again to [Alexander and Imeraj, 2020] [3] and [Woebekking, 2021] [54] for detailed discussions on crypto volatility. Let us denote by $V_{S}(t_{0},t)$ the variance swap on $S$ initiated at time $t_{0}$ and expiring at time $t$. Then, under continuous monitoring, the swap payoff to the long party is nothing but the quadratic variation difference of log-returns cumulated from $t_{0}$ to $t$, with spot price given by
@@
where the notation $\mathsf{E}_{t_{0}}[\cdot]$ means $\mathsf{E}[\cdot|\mathscr{F}_{t_{0}}]$. Clearly, (2.2.2) is well-defined because $S$ in (2.2.1) is a semimartingale.
The following proposition formally expresses the conditional characteristic function of the log-price in terms of the base characteristic functions ($\phi_{X_{1}}$ and $\phi_{Y_{1}}$) and the above variance swap price.
@@
In the above expression, apart from the existence of $\phi_{X_{1}}(-{\rm i})$ and $\phi_{Y_{1}}(-{\rm i}\rho)$ by assumption, note that the integrand $\phi_{Y_{1}}(\rho u-H(t,s)({\rm i}\log\phi_{X_{1}}(u)+u\log\phi_{X_{1}}(-{\rm i})))$ is automatically well-defined provided that the integrated kernel $H$ is nonnegative, which condition is satisfied by all fractional kernels $h$ to be considered. Indeed, in this case one has $(H(t,s)\mathrm{Re}\log\phi_{X_{1}}(u))Y_{1}\leq 0$, $\mathsf{Q}$-a.s., for any $s\leq t$ and $u\in\mathbb{R}$. Also, $H$ is by construction bounded, hence nonsingular. In general, $H$ and its further integral (with respect to $s\in[t_{0},t)$) can both be easily written in closed form, whereas it is difficult to obtain an explicit expression for the $s$-integral involving $\log\phi_{Y_{1}}$, [^9] The reason is that products of power and exponential functions do not produce elementary change of variables under the integral sign; see [Wang and Xia, 2022, Appx. B] [53] for details. except when $h$ takes a piecewise form, to be shown in Section 4.
@@
For any fixed $0\leq t_{0}<t$, the conditional distribution of the log-price, $\mathcal{L}(\log S_{t}|\mathscr{F}_{t_{0}})$, is absolutely continuous with respect to the Lebesgue measure.
We finish this section with a brief discussion of how the crypto price dynamics under the physical measure, say $\mathsf{P}$, equivalent to $\mathsf{Q}$, may be reverse-engineered from (2.2.1). One possibility under which the Lévy properties (namely increment independency and stationarity) of $X$ and $Y$ stay is with some constant $\zeta>0$ and the Radon–Nikodym derivative
$\frac{{\rm d}\mathsf{Q}}{{\rm d}\mathsf{P}}\bigg{|}_{\mathscr{F}_{t}}=\mathcal{E}\bigg{(}\zeta\log\frac{S}{S_{0}}\bigg{)}_{t},\quad t\geq 0,$ (2.2.5)
where $\mathcal{E}$ denotes the Doléans–Dade exponential. This type of measure change, known as the Esscher transform (see, e.g., [Madan and Milne, 1991] [42]), is workable in the present context since the log-cumulative return process $\log(S/S_{0})$ has a finite second moment. [^10] In the language of [Jacod and Shiryaev, 2010] [33], this means that the process $\log(S/S_{0})$ is “exponentially special.” We refer to [Hubalek and Sgarra, 2009] [30] for a profound discussion of this type of measure changes specifically for the Barndorff-Nielsen–Shephard model, which has $X$ being a Brownian motion and $h$ the exponential kernel.
@@
for $(t,z)\in\mathbb{R}_{+}\times(\mathbb{R}\setminus\{0\})$. Then, according to [Jacod and Shiryaev, 2010, Chap. III Thm. 7.18 and Thm. 7.23] [33], if the measure change is achieved through (2.2.5), it suffices that $\log(S/S_{0})$ have the $\mathsf{P}$-semimartingale characteristics
@@
with the understanding that the quantities $\phi_{X_{1}}(-{\rm i})$ and $\phi_{Y_{1}}(-{\rm i}\rho)$ take the same values as in (2.2.1). This implies that the price dynamics under $\mathsf{P}$ can be formulated as
@@
where $\tilde{X}$ and $\tilde{Y}$ are Lévy processes with respective Lévy characteristics $(\mu_{\tilde{X}},\sigma_{\tilde{X}},\nu_{\tilde{X}})$ and $(\mu_{\tilde{Y}},0,\nu_{\tilde{Y}})$, while the business time $\mathcal{T}$ still uses (2.1.1) except with $Y$ replaced by $\tilde{Y}/\rho$ (provided $\rho\neq 0$). In this structure, we stress that the processes $\tilde{X}$ and $\tilde{Y}$ have the same path regularities – measured e.g. by the Blumenthal–Getoor index for the jump components – as $X$ and $\rho Y$, respectively, and the volatility leverage effect is embedded in $\tilde{Y}$, which is nonetheless no longer a $\rho$-scaled Lévy process. Therefore, (2.2.7) also makes it possible to develop estimation methods for the proposed FSV model.

### v3

@@ 1 Introduction
In the present paper we introduce a class of fractional stochastic-volatility models that are capable of reconciling various prevailing stylized facts in financial markets. Broadly speaking, our main objective is to design a general model framework that accommodates both flexible jump dynamics and certain path-dependence properties, and simultaneously achieve fast, semi-analytical valuation. The proposed models address the shortcomings of various existing ones that struggle to provide comprehensive modeling and calibration within a reasonable time frame, while they also maintain parametric parsimony by adding only one or two optional parameters, depending on the desired jump behavior. Following this model framework, we revisit the pricing–hedging problem of derivatives in the cryptocurrency market, where significant price and volatility jumps have been well-documented. Particular attention is drawn to inverse contracts, namely those that are settled in cryptos (    C    ) despite the underlying being valued in a fiat currency (e.g., USD (\$)), which account for the majority of actively traded options across crypto exchanges (including the Deribit exchange).
@@ 1.1 Jumps in crypto prices and volatility
From a derivative pricing perspective, crypto (especially Bitcoin) price dynamics have been explored extensively in recent studies. To mention a few, [Madan et al., 2019] [42] considered calibrating a series of Markov models on Bitcoin options, concluding that models with infinitely active jumps and volatility clustering are generally favorable in terms of pricing errors. This aspect was also strengthened in [Hou et al., 2020] [28], which evidenced the importance of jumps in crypto prices and volatility, as well as their co-movements, by considering both statistical estimation and option pricing under two stochastic-volatility models with correlated jumps from [Duffie et al., 2000] [19] and [Bandi and Renò, 2016] [6], respectively. Significantly, [Hou et al., 2020] [28] has documented an inverse leverage effect, indicating positive serial correlations between crypto price and volatility jumps. Notably, while the diffusive leverage effect may remain negative or insignificant, the inverse leverage effect is particularly pronounced during periods of large fluctuations, which phenomenon is uniquely captured by jump models. We refer to [Huang et al., 2022] [29] for a detailed discussion based on analyzing crypto time series data. Besides, the papers [Jalan et al., 2021] [34] and [Hilliard and Ngo, 2022] [26] also stressed, respectively, the importance of stochastic volatility and price jumps by comparing the pricing performances of a regime-switching stochastic-volatility model and a jump-diffusion model with stochastic convenience yield to that of the Black–Scholes model. Furthermore, [Cao and Celik, 2021] [13] studied the valuation of Bitcoin options under an equilibrium setting with various sources of jump risks, again highlighting the undervaluation of deeply-in-the-money options by the Black–Scholes model.
@@ 2 Model framework
We assume that the crypto market evolves in continuous time $t\geq 0$ and is governed by a stochastic basis $(\Omega,\mathcal{F},\mathsf{P},\mathbb{F}\equiv\{\mathcal{F}_{t}:t\geq 0\})$. The market filtration $\mathbb{F}$ satisfies the usual conditions and contains all market information, and $\mathsf{P}$ is the physical measure. All stochastic processes to appear are understood to be supported on this basis.
For a generic crypto (    C    ), our modeling approach adopts the canonical exponential framework for representing its price dynamics, which is substantially equivalent to specifying a model for the associated log-return process. To this end, we write $S\equiv(S_{t})$ for the crypto price, henceforth the crypto/USD conversion rate, and postulate that it has the following general dynamics under $\mathsf{P}$:
$S_{t}=S_{0}e^{\xi_{t}}\;(\$),\quad t\geq 0,$ (2.1)
where $S_{0}$ is the spot rate and $\xi\equiv(\xi_{t})$ stands for the log-cumulative return, i.e., $\xi=\log(S/S_{0})$. A wide range of well-known existing models can largely be seen as outcomes of specifying a particular stochastic process for $\xi$ – such as pure diffusion processes in the Black–Scholes and Heston models, or processes incorporating jumps, as in jump-diffusion and Gaussian mixture models; compare, e.g., [Madan, 2019, Sect. 3] [42].
Within the general structure (2.1), we introduce the FSV model through a hierarchical framework that nests a base process governing return dynamics and a latent component representing stochastic volatility. These elements will be detailed in the following two subsections in proper order. With our primary focus on valuation, we conduct the analysis directly under a preselected (risk-neutral) pricing measure $\mathsf{Q}$, equivalent to the physical measure $\mathsf{P}$.
@@ 2.1 Fractional stochastic volatility with jumps
When jumps occur in both price and volatility dynamics, a common yet powerful technique that avoids explicitly specifying an instantaneous variance (or squared volatility) process is to incorporate a time-change argument, which works by replacing the calendar time $\imath\equiv(t)$ with a random, constantly increasing business time, denoted as $\mathcal{T}\equiv(\mathcal{T}_{t})$. Intuitively speaking, the business time serves to capture variations in trading activity, incorporate return autocorrelations, and create volatility clustering. Its theoretical validity is deep-rooted in the so-called “theorem of time change” by [Monroe, 1978] [43], which states that any semimartingale can be represented as a time-changed Brownian motion. By extension, the time change of a generally semimartingale is valid provided that the sample paths of $\mathcal{T}$ are continuous $\mathsf{Q}$-a.s. With the monotonicity of $\mathcal{T}$, this implies the typical construction
@@
In general, since volatility is mean-reverting in the long run, a natural model choice for $A$ in the presence of jumps is a Lévy-driven Ornstein–Uhlenbeck process. In the present context, we modify the linear structure with kernel convolution for short-range dependence, by adapting [Wang and Xia, 2022, Sect. 2.2] [52]. Such modification readily leads to fractional Ornstein–Uhlenbeck processes discussed in [Wolpert and Taqqu, 2004] [54], and these processes, having many attractive analytical properties, belong to the broader class of Lévy-driven Volterra processes detailed in [Barndorff-Nielsen et al., 2014] [7].
More precisely, let $Y\equiv(Y_{t})$ be a square-integrable Lévy subordinator (nondeterministic by default), with (Lévy) characteristics $(\mu_{Y},0,\nu_{Y})$ under $\mathsf{Q}$, where $\mu_{Y}\geq 0$ and the Lévy measure $\nu_{Y}$ is non-atomic and corresponds to a Poisson random measure $N_{Y}$ defined on $(\mathbb{R}_{++},\mathbb{R})$, such that $Y_{1}$ admits the characteristic exponent [^7] By default, expectation is taken under the pricing measure $\mathsf{Q}$, i.e., $\mathsf{E}\equiv\mathsf{E}^{\mathsf{Q}}$.
@@
where $\kappa>0$ and $d\in(1/2,1)$ are a mean-reversion parameter and a fraction parameter, respectively. The fractional counterpart of $Y$ is constructed as the following $h$-convoluted stochastic integral:
@@
A notably simple choice of $h$ is the exponential-Riemann–Liouville product kernel (a.k.a. the gamma kernel) (see (4.1.1)), $h(t,s)\equiv h(t-s)=e^{-\kappa(t-s)}(t-s)^{d-1}/\mathrm{\Gamma}(d)$, where $\mathrm{\Gamma}(\cdot)$ is the usual Gamma function and which is widely adopted in the rough volatility literature due to its structural simplicity, including in constructing the rough Heston model (see, e.g., [El Euch and Rosenbaum, 2019] [21]) and fractional Ornstein–Uhlenbeck processes in [Wolpert and Taqqu, 2004] [54]. We will consider other kernel forms as appearing in [Wang and Xia, 2022, Sect. 2.2] [52] later in Section 4.1, and some other special forms of $h$ can be found in [Barndorff-Nielsen et al., 2014, Sect. 3.1] [7].
@@ 2.2 Time-changed processes and crypto prices
Based on (2.1.1), the business time incorporating fractional stochastic volatility is constructed as the time integral of $A$ in (2.1.5), and a major portion of shocks to the crypto price are governed by some $\mathcal{T}$-changed Lévy process, denoted $X\equiv(X_{t})$. Here, $X$ is a real-valued Lévy process with $\mathsf{Q}$-characteristics $(\mu_{X},\sigma_{X},\nu_{X})$, admitting the characteristic exponent
@@
Likewise, we have assumed that $X_{1}$ has finite moments. Also, to capture continual price fluctuations, it is reasonable to assume that either $\sigma_{X}>0$ or $\nu_{X}$ is infinite. The process $X$ is a base process describing return innovations without volatility clustering, which can already have jumps of flexible activity and scales. With stochastic volatility, the crypto price (crypto/USD conversion rate) under the pricing measure $\mathsf{Q}$ is formally given by
@@
where $\rho\in\mathbb{R}$ is a leverage parameter channeling the (direct or inverse) leverage effect between prices and volatility. The existence of the (positive) quantities $\phi_{X_{1}}(-{\rm i})$ and $\phi_{Y_{1}}(-{\rm i}\rho)$ is the same as the requirement that $\mathsf{E}[e^{X_{1}}]<\infty$ and $\mathsf{E}[e^{\rho^{+}Y_{1}}]<\infty$; together, these quantities act as exponential compensators to guarantee the martingale property of $S$ under $\mathsf{Q}$.
The price model (2.2.1) can be seen as a generalized version of the well-known Barndorff-Nielsen–Shephard model ([Barndorff-Nielsen and Shephard, 2001] [8]) in which $X$ is particularly a Brownian motion and $A$ is an ordinary Ornstein–Uhlenbeck process, i.e., one with $h(t,s)=e^{-\kappa(t-s)}$, $s\in[0,t)$. An extension where $X$ is a general Lévy process has been proposed and studied by [Yamazaki, 2016] [59], from which the key difference in the present framework mainly lies in the (singular) kernel function $h$.
To enable dynamic valuation of crypto derivatives, we need the conditional distribution of the log-price $\log S_{t}|\mathcal{F}_{t_{0}}$, for generic time points $0\leq t_{0}<t\leq T$, where $T>0$ is a predetermined expiry date. In the presence of stochastic volatility, dynamic valuation is easily achievable provided access to corresponding volatility derivatives, such as a variance swap which entitles the long party to receive a floating leg equal to the realized variance taken over the swap’s life at the expiry $T>0$ in exchange for a preset fixed payment such that the swap has zero value at the time of issuance. We refer again to [Alexander and Imeraj, 2020] [3] and [Woebekking, 2021] [53] for detailed discussions on crypto volatility. Let us denote by $V_{S}(t_{0},t)$ the variance swap on $S$ initiated at time $t_{0}$ and expiring at time $t$. Then, under continuous monitoring, the swap payoff to the long party is exactly the quadratic variation difference of log-returns cumulated from $t_{0}$ to $t$, with spot price given by
@@
where the notation $\mathsf{E}_{t_{0}}[\cdot]$ means $\mathsf{E}[\cdot|\mathcal{F}_{t_{0}}]$. Clearly, (2.2.2) is well-defined because $S$ in (2.2.1) is a semimartingale.
The following proposition expresses the conditional characteristic function of the log-price in terms of the base characteristic functions ($\phi_{X_{1}}$ and $\phi_{Y_{1}}$) and the above variance swap price.
@@
Technically, in the above expression, apart from the existence of $\phi_{X_{1}}(-{\rm i})$ and $\phi_{Y_{1}}(-{\rm i}\rho)$ by assumption, note that the integrand $\phi_{Y_{1}}(\rho u-H(t,s)({\rm i}\log\phi_{X_{1}}(u)+u\log\phi_{X_{1}}(-{\rm i})))$ is automatically well-defined provided that the integrated kernel $H$ is nonnegative, which condition is satisfied by all fractional kernels $h$ to be considered. Indeed, in this case one has $(H(t,s)\mathrm{Re}\log\phi_{X_{1}}(u))Y_{1}\leq 0$, $\mathsf{Q}$-a.s., for any $s\leq t$ and $u\in\mathbb{R}$. Also, $H$ is by construction bounded, hence nonsingular. In general, $H$ and its further integral (with respect to $s\in[t_{0},t)$) can both be easily written in closed form, whereas it is difficult to obtain an explicit expression for the $s$-integral involving $\log\phi_{Y_{1}}$, [^9] The reason is that products of power and exponential functions do not produce elementary change of variables under the integral sign; see [Wang and Xia, 2022, Appx. B] [52] for details. except when $h$ takes a piecewise form, to be shown in Section 4.
@@
For any fixed $0\leq t_{0}<t$, the conditional distribution of the log-price, $\mathcal{L}(\log S_{t}|\mathcal{F}_{t_{0}})$, is absolutely continuous with respect to the Lebesgue measure.
We finish this section with a brief discussion of how the crypto price dynamics (2.1) under the original physical measure $\mathsf{P}$ may be reverse-engineered from (2.2.1). One possibility under which the Lévy properties (namely increment independency and stationarity) of $X$ and $Y$ stay is with some constant $\zeta>0$ and the Radon–Nikodym derivative
$\frac{{\rm d}\mathsf{Q}}{{\rm d}\mathsf{P}}\bigg{|}_{\mathcal{F}_{t}}=\mathcal{E}\bigg{(}\zeta\log\frac{S}{S_{0}}\bigg{)}_{t},\quad t\geq 0,$ (2.2.5)
where $\mathcal{E}$ denotes the Doléans–Dade exponential. This type of measure change, known as the Esscher transform (see, e.g., [Madan and Milne, 1991] [41]), is workable in the present context since the log-cumulative return process $\xi=\log(S/S_{0})$ has a finite second moment. [^10] In the language of [Jacod and Shiryaev, 2010] [33], this means that the process $\log(S/S_{0})$ is “exponentially special.” We refer to [Hubalek and Sgarra, 2009] [30] for a discussion of this type of measure changes specifically for the Barndorff-Nielsen–Shephard model, which has $X$ being a Brownian motion and $h$ the exponential kernel.
@@
for $(t,z)\in\mathbb{R}_{+}\times(\mathbb{R}\setminus\{0\})$. Then, according to [Jacod and Shiryaev, 2010, Chap. III Thm. 7.18 and Thm. 7.23] [33], if the measure change is achieved through (2.2.5), it suffices that $\xi$ have the $\mathsf{P}$-semimartingale characteristics
@@
with the understanding that the quantities $\phi_{X_{1}}(-{\rm i})$ and $\phi_{Y_{1}}(-{\rm i}\rho)$ take the same values as in (2.2.1). This implies that the price dynamics (2.1) can be formulated as
@@
with $\xi=\tilde{X}_{\mathcal{T}}+\tilde{Y}$. Here, $\tilde{X}$ and $\tilde{Y}$ are Lévy processes with respective Lévy characteristics $(\mu_{\tilde{X}},\sigma_{\tilde{X}},\nu_{\tilde{X}})$ and $(\mu_{\tilde{Y}},0,\nu_{\tilde{Y}})$, while the business time $\mathcal{T}$ still uses (2.1.1) except with $Y$ replaced by $\tilde{Y}/\rho$ (provided $\rho\neq 0$). The processes $\tilde{X}$ and $\tilde{Y}$ have the same path regularities – measured e.g. by the Blumenthal–Getoor index for the jump components – as $X$ and $\rho Y$, respectively, and the volatility leverage effect is embedded in $\tilde{Y}$, which is nonetheless no longer a $\rho$-scaled Lévy process. Therefore, (2.2.7) also makes it possible to develop estimation methods for the proposed FSV model.